OpenStudy (anonymous):

Amanda races her bicycle for 116m . A wheel of her bicycle turns 29 times as the bicycle travels this distance. What is the diameter of the wheel? Use the value 3.14 for Pi . Round your answer to the nearest tenth. Do not round any intermediate steps. Amanda races her bicycle for 116m . A wheel of her bicycle turns 29 times as the bicycle travels this distance. What is the diameter of the wheel? Use the value 3.14 for Pi . Round your answer to the nearest tenth. Do not round any intermediate steps. @Mathematics

OpenStudy (anonymous):

116 = 2*Pi*Radius*29 Radius = 116/(2*Pi*29)

OpenStudy (anonymous):

so would the answer be 1.57

OpenStudy (anonymous):

cause my system says its wrong but thats the one i got

OpenStudy (anonymous):

it is somewhere close to 1.279. try calculating again, you will get it correct

OpenStudy (anonymous):

did you get correct answer?

OpenStudy (phi):

The circumference is 116/29 = 4 m You know C= 2pi r or C= pi d, where d is the diameter so here, we have d= C/pi, d= 4/3.14 = 1.2739 ***rounded the nearest tenth***, we get 1.3